# Exact Correlation Functions for Dual-Unitary Quantum circuits with exceptional points

### v2

@@ Abstract
Dual-unitary quantum circuits can provide analytic spatiotemporal correlation functions of local operators from transfer matrices, enriching our understanding of quantum dynamics with exact solutions. Nevertheless, a full understanding is still lacking as the case of a non-diagonalizable transfer matrix with exceptional points has less been investigated. In this paper, we give an inverse approach for constructing dual-unitary quantum circuits with exceptional points in the transfer matrices, by establishing relations between transfer matrices and local unitary gates. As a consequence of the coalesce of eigenvectors, the correlation functions exhibit a polynomial modified exponential decay, which is significantly different from pure exponential decay, especially at early stages. Moreover, we point out that the Hamiltonian evolution of a kicked XXZ spin chain can be approximately mapped to a dual-unitary circuit with exceptional points by Trotter decomposition. Finally, we investigate the dynamics approaching and at exceptional points, showing that behaviors of correlation functions are distinct by Laplace transformation.
@@ I Introduction
In this paper, we give explicit constructions of dual-unitary quantum circuits whose transfer matrices have EPs and investigate their physical consequences. The construction follows an inverse method, which starts from a transfer matrix with a Jordan block and then derives the local dual-unitary gates, by leveraging constraints of the relations between transfer matrices and the local dual-unitary gates. We find that the transfer matrix with EPs can introduce a polynomial modification on the exponential decay spatiotemporal correlation functions, which stems from the coalesce of eigenvectors. Specifically, we consider both $2\times 2$ and $3\times 3$ Jordan blocks cases of the transfer matrix for qubits. Correspondingly, we propose lattice models of kicked spin XXZ chains whose Hamiltonian evolution at infinite temperature can be approximated by the dual-unitary circuits, due to both of them having similar behaviors of the spatiotemporal correlation functions. Finally, we investigate the dynamics of dual-unitary circuits approaching and at EPs. We find that one can distinguish three different behaviors of the correlation function nearing and at EPs by Laplace transformation.
@@ II Exactly solvable spatiotemporal correlation function in dual-unitary quantum circuit
$\begin{split}U^{\dagger}U=&UU^{\dagger}=\mathbbm{1}\\ \tilde{U}^{\dagger}\tilde{U}=&\tilde{U}\tilde{U}^{\dagger}=\mathbbm{1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1},}\\ \end{split}$ (1)
@@
The classification only considers eigenvalues of the transfer matrix ±$\mathcal{M}_{\pm}(a)$. However, the transfer matrix in general is not Hermitian and may contain Jordan blocks. In other words, the transfer matrix may have EPs. Consequently, the dynamical correlation function should have a polynomial-enhanced behavior. Although this has been noticed in Ref. Bertini et al. (2019), the construction of a dual-unitary quantum circuit where the transfer matrix has EPs has not been explicitly given, leading to a lack of a full classification of dual-unitary quantum circuits.
@@ III Transfer matrix with EPs
Although the parameterization of general dual-unitary two-qudit gates is still not available, the dual-unitary two-qubit gate can be completely parameterized as Bertini et al. (2019),
@@ III.1 Transfer matrix with $2\times 2$ Jordan Block
where $R^{(2)}={\begin{pmatrix}r_{1}&0&l\\ 0&r_{2}&0\\ 0&0&r_{1}\\ \end{pmatrix}}$, $\textbf{0}_{r}={\begin{pmatrix}0&0&0\end{pmatrix}}$, and $\textbf{0}_{c}=\textbf{0}_{r}^{T}$ with $T$ denoting matrix transposition. Here, the parameters $r_{1}$, $r_{2}$, $l$ are related to the specific dual-unitary circuit, the $(2)$ in the upper right corner of $\mathcal{M}$ denotes a $2\times 2$ Jordan Block in the transfer matrix $\mathcal{M}$. To realize a system with the above transfer matrix, one should find the corresponding local evolution operator under the dual-unitary conditions. Without loss of generality, we consider a simple case. The corresponding local dual-unitary evolution operator satisfies the form,
@@ III.2 Transfer matrix with $3\times 3$ Jordan Block
where $R^{(3)}=r\mathbbm{1}_{3}+l_{1}S_{+}+l_{2}S_{+}^{2}$ with $S_{+}={\begin{pmatrix}0&1&0\\ 0&0&1\\ 0&0&0\\ \end{pmatrix}}$ and $\mathbbm{1}_{3}$ is a 3-dimensional identity matrix. Here, the parameters $r$, $l_{1}$, and $l_{2}$ are related to the specific dual-unitary circuit, and the $R^{(3)}$ can be decomposed into a standard $3\times 3$ Jordan block. We keep the parameterization of $R^{(3)}$ instead of a standard Jordan block as it is more convenient for the following derivation. To realize a system with the above transfer matrix, one should find the corresponding local evolution operator under the dual-unitary conditions. Without loss of generality, we consider the corresponding local evolution operator as,
@@ III.3 Kicked XXZ spin chain
$H=H_{XXZ}+\sum_{m=1}^{\infty}\delta\left(\frac{m}{2}\right)(H_{Z}+H_{Y}),$ (32)
where,
@@
Furthermore, we calculate the dynamical correlation function of an infinite temperature state under the kick XXZ model Eq. (32). Interestingly, the light-cone correlation function above can be fitted by the function type of corresponding analytical results Eq. (23) and Eq. (31), as shown in Fig. 3 and Fig. 4. Moreover, the dynamical correlation function is mainly concentrated on the light-cone, which implies that this model can be taken as an approximation of the dual-unitary circuits.
@@
In this section, we first give the analytical results of the correlation function approaching EPs. Then, we analyze behaviors of correlation functions by Fourier transformation and Laplace transformation. It is shown that Laplace transformation can well capture the polynomial modification of the correlation function and thus can distinguish the behaviors approaching and at EPs of the transfer matrix.
@@ IV.1 $2\times 2$ Jordan Block
where $l^{\prime}=\frac{-\cos{(4\Phi-2\delta)}+\frac{1}{2}\sin{2\delta}\sin{4\Phi}}{\cos^{2}{2\Phi}}$. This result exhibits exponential decay rather than polynomial behavior. Moreover, when $\delta<0$, $\Delta$ is a real number. For this case, the correlation function exhibits the superposition of two different modes of exponential decay, which can be observed by Laplace transformation as,
$L_{s}[C^{x,z}_{\delta\neq 0}(t)]=\frac{l^{\prime}}{\Delta}\left[\frac{1}{s-2\log{E_{1}}}-\frac{1}{s-2\log{E_{2}}}\right].\$ (41)
For $\delta=0$, the exceptional points, the Laplace transformation is,
$L_{s}[C^{x,z}(t)]=\frac{2l}{r_{1}(s-2\log{r_{1}})^{2}},$ (42)
@@
which shows exponential decay with one mode. For $\delta>0$, $E_{1}$ and $E_{2}$ turn to complex numbers. Although the Laplace transformation for $C^{x,z}_{\delta>0}$ is the same as the case of $\delta<0$, the result of $L_{s}[C^{x,z}_{\delta>0}]$ is very different from the case of $\delta\leq 0$, as shown in Fig. 5 (g) and (i). Furthermore, due to $E_{1}$ and $E_{2}$ are complex numbers, the correlation function $C^{x,z}_{\delta>0}$ will show oscillation behavior. By introducing Fourier transformation, we obtain the frequency function,
$f_{\delta>0}(\omega)=\frac{l^{\prime}4i\theta_{\delta}}{\Delta(4\theta_{\delta}^{2}-\omega^{2}+4\log{|E_{1}|}(\log{|E_{1}|}-i\omega))},$ (43)
where $\theta_{\delta}=\arctan{\frac{|\Delta|}{E_{1}+E_{2}}}$. By the way, we calculate the Fourier transformation of the correlation function $C^{x,z}$ as,
$\begin{split}f(\omega)=-\frac{2l}{r_{1}(\omega+2i\log{r_{1}})^{2}},\end{split}$ (44)
for $\delta=0$, and,
$\begin{split}f_{\delta<0}(\omega)=\frac{l^{\prime}}{\Delta}\frac{\log{E_{1}}-\log{E_{2}}}{(-\log{E_{1}}+i\omega)(-\log{E_{2}}+i\omega)},\end{split}$ (45)
for $\delta<0$.
@@
As shown in Fig. 5(a), (b), and (c), the properties of the correlation function $C^{x,z}$ are different for approaching and at EPs. For $\delta<0$, the correlation function $C^{x,z}$ shows exponential behavior with two modes, which include one mode decay and one mode growth. These two modes can be clearly depicted by Laplace transformation with two points of divergence, as shown in Fig. 5 (g). Moreover, this behavior is fundamentally different from the case of $\delta=0$, which exhibits a polynomial enhancement with one mode exponential decay. Although Fig. 5 (a) and (b) show similar behavior, increasing firstly and then decreasing, the Laplace transformation of case $\delta=0$ have only one point of divergence, as shown in Fig. 5 (h). However, for $\delta>0$, the correlation function $C^{x,z}$ shows oscillation with exponential decay [as shown in Fig. 5 (c)], which is significantly different from the previous two cases. On the other hand, as shown in Fig. 5 (d), (e) and, (f), Fourier analysis gives signals that are not that clear compared to Laplace transformation.
@@
As a consequence, the rich properties of the correlation functions $C^{x,z}$ for approaching and at EPs can be clearly observed by Laplace transformation, which is due to the differences of the transfer matrices. For $\delta=0$, the coalesce of eigenvectors results in the correlation showing polynomial behavior. For $\delta<0$, the correlation shows two modes of exponential decay, because the transfer matrix can be diagonalized as a real spectrum. While, for $\delta>0$, due to the complex spectrum of the transfer matrix, the correlation function exhibits the oscillation behavior with one mode of exponential decay.
@@
Figure 5: The correlation function, Fourier amplitude, and Laplace transformation of correlation function $C^{x,z}_{\delta}$ for transfer matrix with $2\times 2$ Jordan block case. (a), (b), and (c) is the correlation function $C^{x,z}_{\delta}$ with different parameters $\delta$. (d), (e), and (f) is the corresponding Fourier amplitude of correlation function $C^{x,z}_{\delta}$. (g), (h), and (i) is the corresponding Laplace transformation of the correlation function $C^{x,z}_{\delta}$.
@@ IV.2 $3\times 3$ Jordan Block
$\begin{split}A_{1}=&\frac{2l_{3}[l_{2}l_{4}+l_{1}(r-r^{\prime}_{1})]+(r^{\prime}_{1}-r^{\prime}_{2}+\Delta^{\prime})[l_{1}l_{3}+l_{2}(r-r^{\prime}_{2})]}{2\Delta^{\prime}[l_{3}l_{4}+(r^{\prime}_{1}-r)(r-r^{\prime}_{2})]}\\ A_{2}=&-\frac{2l_{3}[l_{2}l_{4}+l_{1}(r-r^{\prime}_{1})]+(r^{\prime}_{1}-r^{\prime}_{2}-\Delta^{\prime})[l_{1}l_{3}+l_{2}(r-r^{\prime}_{2})]}{2\Delta^{\prime}[l_{3}l_{4}+(r^{\prime}_{1}-r)(r-r^{\prime}_{2})]}\\ A_{3}=&-\frac{l_{1}l_{3}+l_{2}(r-r^{\prime}_{2})}{l_{3}l_{4}+(r^{\prime}_{1}-r)(r-r^{\prime}_{2})}.\\ \end{split}$ (50)
@@
These results imply that the correlation function will exhibit multiple modes of exponential decay rather than polynomial behavior. For $\delta<0$, $\Delta^{\prime}$ is a real number. In this case, the correlation function $C^{x,z}$ is a superposition of three different exponential decay modes, which can be described by Laplace transformation as,
$L_{s}[C^{x,z}_{\delta\neq 0}(t)]=\frac{A_{1}}{s-2\log{E_{1}}}+\frac{A_{2}}{s-2\log{E_{2}}}+\frac{A_{3}}{s-2\log{E_{3}}}.\$ (51)
For $\delta=0$, the Laplace transformation of correlation function $C^{x,z}$ is,
$L_{s}[C^{x,z}(t)]=\frac{4l_{1}^{2}}{r^{2}(s-2\log{r})^{3}}+\frac{2l_{2}r-l_{1}^{2}}{r^{2}(s-2\log{r})^{2}},$ (52)
@@
which shows that the correlation function exhibits exponential decay with one mode. However, for $\delta>0$, $\Delta^{\prime}$ is an image number, which implies that the correlation function will exhibit oscillation behavior. By introducing the Fourier transformation, we obtain the function of frequency as,
$\begin{split}f_{\delta>0}(\omega)=&\frac{iA_{1}}{2\theta_{\delta}-\omega-i2\log{|E_{1}|}}-\frac{iA_{2}}{2\theta_{\delta}+\omega+i2\log{|E_{1}|}}\\ &-\frac{iA_{3}}{\omega+i2\log{E_{3}}},\\ \end{split}$ (53)
where $\theta_{\delta}=\arctan{\frac{|\Delta^{\prime}|}{E_{1}+E_{2}}}$. Furthermore, we calculate the Fourier transformation of the correlation function $C^{x,z}$ as,
$f(\omega)=\frac{l_{1}^{2}-2l_{2}r}{r^{2}(\omega+i2\log{r})^{2}}+\frac{i4l_{1}^{2}}{r^{2}(\omega+i2\log{r})^{3}}$ (54)
for $\delta=0$, and,
$f_{\delta<0}(\omega)=\frac{A_{1}}{i\omega-2\log{E_{1}}}+\frac{A_{2}}{i\omega-2\log{E_{2}}}+\frac{A_{3}}{i\omega-2\log{E_{3}}}.$ (55)
for $\delta<0$.
@@
Although the correlation functions look very similar [shown in Fig. 6 (a), (b), and (c)], their dynamical properties are fundamentally different. Specifically, for $\delta<0$, the correlation function shows exponential decay behavior with three modes, which is clearly described by Laplace transformation with three divergent points, as shown in Fig. 6 (g). However, for $\delta>0$, the Laplace transformation shows only one divergent point, because two eigenvalues of the transfer matrix turn into complex numbers, which will make the correlation function show oscillating rather than exponential decay behaviors. Moreover, this is different from the case of $\delta=0$, the polynomial-enhanced with exponential decay, whose Laplace transformation appears as a higher order divergence point, as shown in Fig. 6 (h) and (i). In contrast, we find that the Fourier transformation can not clearly distinguish the different correlation functions with different dynamical properties for approaching and at EPs.
@@
Figure 6: The correlation function, Fourier amplitude, and Laplace transformation of correlation function $C^{x,z}_{\delta}$ for transfer matrix with $3\times 3$ case. (a), (b), and (c) is the correlation function $C^{x,z}_{\delta}$ with different parameters $\delta$. (d), (e), and (f) is the corresponding Fourier amplitude of correlation function $C^{x,z}_{\delta}$. (g), (h), and (i) is the corresponding Laplace transformation of the correlation function $C^{x,z}_{\delta}$.
@@
In summary, we have constructed the local evolution gates, whose transfer matrix includes $2\times 2$ and $3\times 3$ Jordan blocks for $d=2$ case. By solving the relationship between the transfer matrix and spatiotemporal correlation function, we have derived the constrained relationship of the parameters of the local evolution operator to the transfer matrix with $2\times 2$ and $3\times 3$ Jordan blocks for $d=2$ case. Moreover, we have proposed the Hamiltonian evolution of a kick XXZ mode as an approximation of the dual-unitary circuits, due to both of them having similar behaviors of the spatiotemporal correlation functions. Lastly, we have shown that distinct behaviors of correlation functions approaching and at EPs can be obtained by Laplace transformation.

### v3

@@ Abstract
Dual-unitary quantum circuits can provide analytic spatiotemporal correlation functions of local operators from transfer matrices, enriching our understanding of quantum dynamics with exact solutions. Nevertheless, a full understanding is still lacking as the case of a non-diagonalizable transfer matrix with exceptional points has hardly been explored. In this paper, we give an inverse approach for constructing dual-unitary quantum circuits with exceptional points in the transfer matrices, by establishing relations between transfer matrices and local unitaries. As a consequence of the coalesce of eigenvectors, the correlation functions exhibit a polynomial modified exponential decay, which is significantly different from pure exponential decay, especially at early stages. Moreover, we point out that the Hamiltonian evolution of a kicked XXZ spin chain can be approximately mapped to a dual-unitary circuit with exceptional points by Trotter decomposition. Finally, we investigate the dynamics approaching and at exceptional points, showing that behaviors of correlation functions are distinct by $Z$ transformation.
@@ I Introduction
In this paper, we give explicit constructions of dual-unitary quantum circuits whose transfer matrices have EPs and investigate their physical consequences. The construction follows an inverse method, which starts from a transfer matrix with a Jordan block and then derives the local dual-unitary gates, by leveraging constraints of the relations between transfer matrices and the local dual-unitary gates. We find that the transfer matrix with EPs can introduce a polynomial modification on the exponential decay spatiotemporal correlation functions, which stems from the coalesce of eigenvectors. Specifically, we consider both $2\times 2$ and $3\times 3$ Jordan blocks cases of the transfer matrix for qubits. Correspondingly, we propose lattice models of kicked spin XXZ chains whose Hamiltonian evolution at infinite temperature can be approximated by the dual-unitary circuits, due to both of them having similar behaviors of the spatiotemporal correlation functions. Finally, we investigate the dynamics of dual-unitary circuits approaching and at EPs. We find that one can distinguish three different behaviors of the correlation function nearing and at EPs by $Z$ transformation.
@@ II Exactly solvable spatiotemporal correlation function in dual-unitary quantum circuit
$\begin{split}U^{\dagger}U=&UU^{\dagger}=\mathbbm{1}\\ \tilde{U}^{\dagger}\tilde{U}=&\tilde{U}\tilde{U}^{\dagger}=\mathbbm{1},\\ \end{split}$ (1)
@@
Here, $\lambda_{i}$ is the eigenvalue of the transfer matrix $\mathcal{M}$ with $i=0,1,2,3$.
The classification only considers eigenvalues of the transfer matrix ±$\mathcal{M}_{\pm}(a)$. However, the transfer matrix in general is not Hermitian and may contain Jordan blocks. In other words, the transfer matrix may have EPs. Consequently, the dynamical correlation function should have a polynomial-enhanced behavior. Although this has been noticed in Ref. Bertini et al. (2019), the construction of a dual-unitary quantum circuit where the transfer matrix has EPs has not been explicitly given, leading to a lack of a full classification of dual-unitary quantum circuits. In additional, it is noted that consequences of non-diagonalizable transfer matrix also show up for quantum transport properties Saha et al. (2023a, b, c), although not in the context of dual-unitary quantum circuit.
@@ III Transfer matrix with EPs
Although the parameterization of general dual-unitary gates are still not available, the dual-unitary two-qubit gate can be completely parameterized as Bertini et al. (2019),
@@ III.1 Transfer matrix with $2\times 2$ Jordan Block
where $R^{(2)}={\begin{pmatrix}r_{1}&0&l\\ 0&r_{2}&0\\ 0&0&r_{1}\\ \end{pmatrix}}$, $\textbf{0}_{r}={\begin{pmatrix}0&0&0\end{pmatrix}}$, and $\textbf{0}_{c}=\textbf{0}_{r}^{T}$ with $T$ denoting matrix transposition. Here, the parameters $r_{1}$, $r_{2}$, $l$ are related to the specific dual-unitary circuit, the $(2)$ in the upper right corner of $\mathcal{M}$ denotes a $2\times 2$ Jordan Block in the transfer matrix $\mathcal{M}$. To realize a system with the above transfer matrix, one should find the corresponding local evolution operator under the dual-unitary conditions. For simplicity, we consider the corresponding local dual-unitary evolution operator which satisfies the form,
@@ III.2 Transfer matrix with $3\times 3$ Jordan Block
where $R^{(3)}=r\mathbbm{1}_{3}+l_{1}S_{+}+l_{2}S_{+}^{2}$ with $S_{+}={\begin{pmatrix}0&1&0\\ 0&0&1\\ 0&0&0\\ \end{pmatrix}}$ and $\mathbbm{1}_{3}$ is a 3-dimensional identity matrix. Here, the parameters $r$, $l_{1}$, and $l_{2}$ are related to the specific dual-unitary circuit, and the $R^{(3)}$ can be decomposed into a standard $3\times 3$ Jordan block. We keep the parameterization of $R^{(3)}$ instead of a standard Jordan block as it is more convenient for the following derivation. To realize a system with the above transfer matrix, one should find the corresponding local evolution operator under the dual-unitary conditions. For simplicity, we consider the corresponding local evolution operator as,
@@ III.3 Kicked XXZ spin chain
$H=H_{XXZ}+\sum_{m=1}^{\infty}\delta\left(t-\frac{m}{2}\right)(H_{Z}+H_{Y}),$ (32)
where $\delta(t-\frac{m}{2})$ is Dirac delta function and,
@@
Furthermore, we calculate the dynamical correlation function of an infinite temperature state under the kick XXZ model Eq. (32). Interestingly, the light-cone correlation function above can be fitted by the function type of corresponding analytical results Eq. (23) and Eq. (31), as shown in Fig. 3 and Fig. 4. Moreover, the dynamical correlation function is mainly concentrated on the light-cone and the light-cone correlation function can be fitting the polynomial modified exponential decay, which implies that this model can be an approximation of the dual-unitary circuits.
@@
In this section, we first give the analytical results of the correlation function approaching EPs. Then, we analyze behaviors of correlation functions by discrete Fourier transformation and $Z$ transformation. It is shown that $Z$ transformation can well capture the polynomial modification of the correlation function and thus can distinguish the behaviors approaching and at EPs of the transfer matrix.
@@ IV.1 $2\times 2$ Jordan Block
where $l^{\prime}=\frac{-\cos{(4\Phi-2\delta)}+\frac{1}{2}\sin{2\delta}\sin{4\Phi}}{\cos^{2}{2\Phi}}$. This result exhibits exponential decay rather than polynomial behavior. Moreover, when $\delta<0$, $\Delta$ is a real number. For this case, the correlation function exhibits the superposition of two different modes of exponential decay, which can be observed by $Z$ transformation as,
$\mathcal{Z}_{\mathbf{z}}[C^{x,z}_{\delta\neq 0}(t)]=\frac{l^{\prime}\mathbf{z}}{\Delta}\left[\frac{1}{\mathbf{z}-{E_{1}}^{2}}-\frac{1}{\mathbf{z}-{E_{2}}^{2}}\right].\$ (41)
For $\delta=0$, the exceptional points, the $Z$ transformation is,
$\mathcal{Z}_{\mathbf{z}}[C^{x,z}(t)]=\frac{2lr_{1}\mathbf{z}}{(\mathbf{z}-{r_{1}}^{2})^{2}},$ (42)
@@
which shows exponential decay with one mode. For $\delta>0$, $E_{1}$ and $E_{2}$ turn to complex numbers. Although the $Z$ transformation for $C^{x,z}_{\delta>0}$ is the same as the case of $\delta<0$, the result of $\mathcal{Z}_{\mathbf{z}}[C^{x,z}_{\delta>0}]$ is very different from the case of $\delta\leq 0$, as shown in Fig. 5 (g) and (i). Furthermore, due to $E_{1}$ and $E_{2}$ are complex numbers, the correlation function $C^{x,z}_{\delta>0}$ will show oscillation behavior. By introducing discrete Fourier transformation, we obtain the frequency function,
$f_{\delta\neq 0}(\omega)=\frac{l^{\prime}}{\Delta}\left[\frac{1}{1-{E_{1}}^{2}e^{i2\omega}}-\frac{1}{1-{E_{2}}^{2}e^{i2\omega}}\right].$ (43)
By the way, we calculate the discrete Fourier transformation of the correlation function $C^{x,z}$ as,
$\begin{split}f(\omega)=\frac{2lr_{1}e^{-i2\omega}}{(e^{-i2\omega}-{r_{1}}^{2})^{2}},\end{split}$ (44)
for $\delta=0$.
@@
As shown in Fig. 5(a), (b), and (c), the properties of the correlation function $C^{x,z}$ are different for approaching and at EPs. For $\delta<0$, the correlation function $C^{x,z}$ shows exponential behavior with two modes, which include one mode decay and one mode growth. These two modes can be clearly depicted by $Z$ transformation with two points of divergence, as shown in Fig. 5 (g). Moreover, this behavior is fundamentally different from the case of $\delta=0$, which exhibits a polynomial enhancement with one mode exponential decay. Although Fig. 5 (a) and (b) show similar behavior, increasing firstly and then decreasing, the $Z$ transformation of case $\delta=0$ have only one point of divergence, as shown in Fig. 5 (h). However, for $\delta>0$, the correlation function $C^{x,z}$ shows oscillation with exponential decay [as shown in Fig. 5 (c)], which is significantly different from the previous two cases. On the other hand, as shown in Fig. 5 (d), (e) and, (f), discrete Fourier analysis gives signals that are not that clear compared to $Z$ transformation.
@@
As a consequence, the rich properties of the correlation functions $C^{x,z}$ for approaching and at EPs can be clearly observed by $Z$ transformation, which is due to the differences of the transfer matrices. For $\delta=0$, the coalesce of eigenvectors results in the correlation showing polynomial behavior. For $\delta<0$, the correlation shows two modes of exponential decay, because the transfer matrix can be diagonalized as a real spectrum. While, for $\delta>0$, due to the complex spectrum of the transfer matrix, the correlation function exhibits the oscillation behavior with one mode of exponential decay.
@@
Figure 5: The correlation function, Fourier amplitude, and $Z$ transformation of correlation function $C^{x,z}_{\delta}$ for transfer matrix with $2\times 2$ Jordan block case. (a), (b), and (c) is the correlation function $C^{x,z}_{\delta}$ with different parameters $\delta$. (d), (e), and (f) is the corresponding Fourier amplitude of correlation function $C^{x,z}_{\delta}$. (g), (h), and (i) is the corresponding $Z$ transformation of the correlation function $C^{x,z}_{\delta}$.
@@ IV.2 $3\times 3$ Jordan Block
$\begin{split}A_{1}=&\frac{2l_{3}[l_{2}l_{4}+l_{1}(r-r^{\prime}_{1})]+(r^{\prime}_{1}-r^{\prime}_{2}+\Delta^{\prime})[l_{1}l_{3}+l_{2}(r-r^{\prime}_{2})]}{2\Delta^{\prime}[l_{3}l_{4}+(r^{\prime}_{1}-r)(r-r^{\prime}_{2})]}\\ A_{2}=&\frac{2l_{3}[l_{2}l_{4}+l_{1}(r-r^{\prime}_{1})]+(r^{\prime}_{1}-r^{\prime}_{2}-\Delta^{\prime})[l_{1}l_{3}+l_{2}(r-r^{\prime}_{2})]}{-2\Delta^{\prime}[l_{3}l_{4}+(r^{\prime}_{1}-r)(r-r^{\prime}_{2})]}\\ A_{3}=&-\frac{l_{1}l_{3}+l_{2}(r-r^{\prime}_{2})}{l_{3}l_{4}+(r^{\prime}_{1}-r)(r-r^{\prime}_{2})}.\\ \end{split}$ (49)
@@
These results imply that the correlation function will exhibit multiple modes of exponential decay rather than polynomial behavior. For $\delta<0$, $\Delta^{\prime}$ is a real number. In this case, the correlation function $C^{x,z}$ is a superposition of three different exponential decay modes, which can be described by $Z$ transformation as,
$\mathcal{Z}_{\mathbf{z}}[C^{x,z}_{\delta\neq 0}(t)]=\mathbf{z}\left(\frac{A_{1}}{\mathbf{z}-{E_{1}}^{2}}+\frac{A_{2}}{\mathbf{z}-{E_{2}}^{2}}+\frac{A_{3}}{\mathbf{z}-{E_{3}}^{2}}\right).\$ (50)
For $\delta=0$, the $Z$ transformation of correlation function $C^{x,z}$ is,
$\mathcal{Z}_{\mathbf{z}}[C^{x,z}(t)]=\frac{2l_{2}r\mathbf{z}}{(\mathbf{z}-r^{2})^{2}}+\frac{{l_{1}}^{2}(3r^{2}+\mathbf{z})}{(\mathbf{z}-r^{2})^{3}},$ (51)
@@
which shows that the correlation function exhibits exponential decay with one mode. However, for $\delta>0$, $\Delta^{\prime}$ is an image number, which implies that the correlation function will exhibit oscillation behavior. By introducing the discrete Fourier transformation, we obtain the function of frequency as,
$f_{\delta\neq 0}(\omega)=\frac{A_{1}}{1-{E_{1}}^{2}e^{i2\omega}}+\frac{A_{2}}{1-{E_{2}}^{2}e^{i2\omega}}+\frac{A_{3}}{1-{E_{3}}^{2}e^{i2\omega}}.$ (52)
Furthermore, we calculate the discrete Fourier transformation of the correlation function $C^{x,z}$ as,
$f(\omega)=\frac{2l_{2}re^{-i2\omega}}{(e^{-i2\omega}-r^{2})^{2}}+\frac{{l_{1}}^{2}(3r^{2}+e^{-i2\omega})}{(e^{-i2\omega}-r^{2})^{3}}$ (53)
for $\delta=0$.
@@
Although the correlation functions look very similar [shown in Fig. 6 (a), (b), and (c)], their dynamical properties are fundamentally different. Specifically, for $\delta<0$, the correlation function shows exponential decay behavior with three modes, which is clearly described by $Z$ transformation with three divergent points, as shown in Fig. 6 (g). However, for $\delta>0$, the $Z$ transformation shows only one divergent point, because two eigenvalues of the transfer matrix turn into complex numbers, which will make the correlation function show oscillating rather than exponential decay behaviors. Moreover, this is different from the case of $\delta=0$, the polynomial-enhanced with exponential decay, whose $Z$ transformation appears as a higher order divergence point, as shown in Fig. 6 (h) and (i). In contrast, we find that the discrete Fourier transformation can not clearly distinguish the different correlation functions with different dynamical properties for approaching and at EPs.
@@
Figure 6: The correlation function, Fourier amplitude, and $Z$ transformation of correlation function $C^{x,z}_{\delta}$ for transfer matrix with $3\times 3$ case. (a), (b), and (c) is the correlation function $C^{x,z}_{\delta}$ with different parameters $\delta$. (d), (e), and (f) is the corresponding Fourier amplitude of correlation function $C^{x,z}_{\delta}$. (g), (h), and (i) is the corresponding $Z$ transformation of the correlation function $C^{x,z}_{\delta}$.
@@
In summary, we have constructed the local evolution gates, whose transfer matrix includes $2\times 2$ and $3\times 3$ Jordan blocks for $d=2$ case. By solving the relationship between the transfer matrix and spatiotemporal correlation function, we have derived the constrained relationship of the parameters of the local evolution operator to the transfer matrix with $2\times 2$ and $3\times 3$ Jordan blocks for $d=2$ case. Moreover, we have proposed the Hamiltonian evolution of a kick XXZ mode as an approximation of the dual-unitary circuits, due to both of them having similar behaviors of the spatiotemporal correlation functions. Lastly, we have shown that distinct behaviors of correlation functions approaching and at EPs can be obtained by $Z$ transformation.